2. A monopolist has a cost function given by TC 250+q+.004q. The inverse market demand for boxes is given by p 8-.0010. The monopolist is curranty able to exclude rivals from the market becaus of a spocial governmental zoning rule (a) What is its output and what price does it charge for boxes? (b) Calculate the firms profit at this output level. (c) Calculate the firms producers surplus at this output level. (d) Calculate the consumers surplus in this situation. 3. The zoning rule is revoked, and the monopolist in problem 2 can no longer exclude others from using the same technology and producing boxes, so the market structure changes from monopoly to perfect competition. (That is, assume that all firms are price-takers and that they produce at minimum average cost in equilibrium.) (a) What will the market price and quantity be? (b) Calculate consumers surplus under this market structure. (c) Calculate aggregate producer profits and producers surplus. surplus (sum of producers and consumers surplus)? 4. The elasticity of demand in the local hardware industry is -1.5, while in the video market it is -4. Which industr has a higher markup over marginal cost (as a percentage of price)? In answering, calculate the markup for each.
(c) Calculate aggregate producer profits and producers surplus (d) Comparing the two market structures for this industry, which has a larger social surplus (sum of producers and consumers surplus)? 4. The elasticity of demand in the local hardware industry is -1.5, while in the video market it is -4. Which industry has a higher markup over marginal cost (as a percentage of price)? In answering, calculate the markup for each. 5. The local bottled water supplier sells water for S0.50 per gallon, while marginal cost is S0.20 per gallon. What is this firms degree of monopoly power, as measured by the Lerner index? 6. What is a network externality? And why might goods that exhibit this externality be more likely to be provided b monopoly suppliers?
Answer #1

2)

Total Cost (TC) = 250+ q +0.004q2

Demand: p = 8 - 0.001Q

a) The monopolist will produce where the marginal revenue equals the marginal cost.

MC = dTC/dq

MC = 1+0.008q

TR = P*Q

TR = 8Q – 0.001Q2

Marginal Revenue(MR) = dTR/dQ

MR = 8-0.002Q

Therefore,

1+0.008q = 8 – 0.002q

0.01q = 7

q = 700

Price = 8 – 0.001*700

Price = 7.3

b) Profit = TR – TC

Profit = 7.3*700 – (250+ 700+1960)

Profit = 5110 - 2910 = $2200

c) Producer surplus = Profit = $2200

d) Consumer Surplus = 0.5*(8 – 7.3)*700 = $245

*We are supposed to do only one question.